5. 14 marks In this problem we will explore the factors of expression
a. 12 marks] Suppose a friend comes to you claiming that "(s ) divides sWrite your friend's claim as a theorem, specifying the domains for each of the variables and using appropriate quantifiers. b. 5 marks Use mathematical induction to prove the theorem! Hint: sk-can be rewritten in the form (s-11) + z where r is an expression involving s, t, k and some constants tk-1) +1 where x is an expression involving s, t, k and some constants n the form SS
Answer #1

We are presenting a consolidated answer here:


Let the given statement be P(n). Then we need to prove that:

P(k): (sn - tn) is divisible by (s - t).

When n = 1, the given statement becomes: (s1 - t1) is divisible by (s - t), which stands true, hence P(1) is true.

Let p(k) be true. Then P(k): sk - tk is divisible by (s-t).

Now, sk + 1 - tk + 1 = sk + 1 - skt + skt - tk + 1 [on adding and subtracting skt]

= sk(s - t) + t(sk - tk), which is divisible by (s - t) [using (i)]

⇒ P(k + 1): sk + 1 - tk + 1is divisible by (s - t)

⇒ P(k + 1) is true, whenever P(k) is true.

Thus as P(1) and P(k + 1) stand true, whenever P(k) is true. Hence, by the Principal of Mathematical Induction, P(n) is true for all n ∈ N.